Work in Process Account Data for Two Months; Cost of Production Reports Pittsburgh Aluminum Company uses a process cost systeACCOUNT Work in Process-Rolling Department ACCOUNT NO. BALANCE DATE ITEM POST. REF. DEBIT CREDIT DEBIT CREDIT 7,512 Sept. 1 SPittsburgh Aluminum Company Cost of Production Report-Rolling Department For the Month Ended September 30 Whole Units EquivalCosts Costs Direct Materials Conversion Total Costs Cost per equivalent unit: Total costs for September in Rolling Department2. Provide the same information for October by recording the October transactions in the four-column work in process account.Pittsburgh Aluminum Company Cost of Production Report-Rolling Department For the Month Ended October 31 Whole Units EquivalenCosts Costs Direct Materials Conversion Total Costs Cost per equivalent unit: Total costs for October in Rolling Department T
Answer #1

Part 1

PITTSBURGH ALUMINUM COMPANY

Cost of Production Report—Rolling Department

For the Month Ended September 30

UNITS

Equivalent Units

Whole Unit

Direct Materials

Conversion

Units charged to production:

Inventory in process, September 1

1200

Received from Smelting Department

27240

Total units accounted for by the Rolling Department

28440

Units to be assigned cost:

Inventory in process, September 1 (60% completed)

1200

0

480

Started and completed in September

25140

25140

25140

Transferred to finished goods in September

26340

25140

25620

Inventory in process, September 30 (20% complete)

2100

2100

420

Total units to be assigned cost

28440

27240

26040

COSTS

Costs

Direct Materials

Conversion

Total

Costs per equivalent unit:

Total costs for September in Rolling Department

138924

57288

Total equivalent units

27240

26040

Cost per equivalent unit

5.10

2.20

Costs assigned to production:

Inventory in process, September 1

7512

Costs incurred in September

196212

Total costs accounted for by the Rolling Department

203724

Cost allocated to completed and partially completed units:

Inventory in process, September 1 balance

7512

To complete inventory in process, September 1

0

1056

1056

Cost of completed September 1 work in process

8568

Started and completed in September

128214

55308

183522

Transferred to finished goods in September

192090

Inventory in process, September 30

9690

924

11634

Total costs assigned by the Rolling Department

203724

37237+20051 =57288

480*2.20= 1056

25140*5.10 = 128214

25140*2.2 = 55308

2100*4.70 = 9690

420*2.20= 924

Part 2

Date

Item

Debit

Credit

Debit

Credit

October 1

Balance

11634

October 31

Smelting Dept 31300 units at $5.3

165890

177524

October 31

Direct labor

48830

226354

October 31

Factory overhead

26288

252642

October 31

Finished goods

October 31

Bal., 1600 units, 80% completed

PITTSBURGH ALUMINUM COMPANY

Cost of Production Report—Rolling Department

For the Month Ended October 31

UNITS

Equivalent Units

Whole Unit

Direct Materials

Conversion

Units charged to production:

Inventory in process, October 1

2100

Received from Smelting Department

31300

Total units accounted for by the Rolling Department

33400

Units to be assigned cost:

Inventory in process, October 1 (20% completed)

2100

0

1680

Started and completed in September

29700

29700

29700

Transferred to finished goods in September

31800

29700

31380

Inventory in process, October 30 (80% complete)

1600

1600

1280

Total units to be assigned cost

33400

31300

32660

COSTS

Costs

Direct Materials

Conversion

Total

Costs per equivalent unit:

Total costs for October in Rolling Department

165890

75118

Total equivalent units

31300

32660

Cost per equivalent unit

5.30

2.30

Costs assigned to production:

Inventory in process, October 1

11634

Costs incurred in October

241008

Total costs accounted for by the Rolling Department

252642

Cost allocated to completed and partially completed units:

Inventory in process, October 1 balance

11634

To complete inventory in process, October 1

0

3864

3864

Cost of completed October 1 work in process

15498

Started and completed in September

157410

68310

225720

Transferred to finished goods in September

241218

Inventory in process, September 30

8480

2944

11424

Total costs assigned by the Rolling Department

252642

48830+26288 = 75118

1680*2.30 = 3864

29700*5.30 = 157410

29700*2.30 = 68310

1600*5.30 = 8480

1280*2.30 = 2944

Part 3

The cost per equivalent unit for direct materials increased from August to October. The cost per equivalent unit for conversion costs increased August to October. These charges should be investigated for their underlying causes, and any necessary corrective actions should be taken.